OpenStudy (anonymous):

Solve this problem on paper using all four steps. A girl scout troop sold cookies. If the girls sold 5 more boxes the second week than they did the first, and if they doubled the sales of the second week for the third week to sell a total of 431 boxes of cookies, how many did they sell each week? They sold___boxes the first week,___boxes the second week, and___boxes the third week.

OpenStudy (sweetburger):

x = first week y = x+5 = second week z = 2(x+5) = third week x+y+z = 431 therefore we substitue x + (x+5) + 2(x+5) = 431 solve for x and then plug that value back into the 3 orginial equations above to see how many boxes of cookies were sold each week

OpenStudy (sweetburger):

Gaben would approve these kind of sales

OpenStudy (anonymous):

thanks for the explanation :) and what sales?

OpenStudy (anonymous):

besides something that haves that number i cant think of

OpenStudy (sweetburger):

ahh nm was a dumb reference

OpenStudy (anonymous):

well sorry if i diddnt catch it

OpenStudy (anonymous):

ugh no

OpenStudy (sweetburger):

alrighty lord gaben take care